OpenStudy (anonymous):

Identify the solutions of 3x^2 + 14x + 8 = 0

OpenStudy (anonymous):

Do you know what the quadratic formula is?

OpenStudy (anonymous):

Or, do you know how to factor?

OpenStudy (anonymous):

x=-+ or -sqrt b2-4ac ----------------------- 2a ?

OpenStudy (anonymous):

Something like that.

OpenStudy (anonymous):

opps i ment x =b+

OpenStudy (anonymous):

BTW, here is the factored form: (x+4)(3x+2)=0

OpenStudy (anonymous):

Do you think you can solve it now?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

(x+4)=0 (3x+2)=0 Now just solve each equation.

OpenStudy (anonymous):

alright thanks so much!

OpenStudy (anonymous):

Yep